OpenStudy (anonymous):

What makes a triangle a “special” right triangle? How can special right triangles help me find side lengths?

OpenStudy (anonymous):

look up formual chart on google scroll to where it say special right triangle and there is a formual you need to use on each side of your triangle

OpenStudy (anonymous):

@Luis_Rivera lol nice i was about to get to that lol

OpenStudy (anonymous):

one of the angles is 90 degree and the other is 45 degree. the sum of three angles of triangle adds to a80 degrees, which means the remaining angle is also 45 degree. S0, if the two angles of a triangle are equal then its isosceles triangle, which means the two sides are also equal. so the other side is also 2 square root of 2. Now to find the hypotenuse: h^2 = p^2 + b^2 (pythagoras theorem) so, h^2 = (2 square root of 2)^2 + (2 square root of 2)^2 = 4x2 + 4x2 = 8+ 8 = 16 Therefore hypotenus = 4 (square root of 16)

OpenStudy (anonymous):

A special right triangle is a right triangle with some regular feature that makes calculations on the triangle easier, or for which simple formulas exist. Read from here: http://en.wikipedia.org/wiki/Special_right_triangles#Right_triangle_whose_angles_are_in_a_geometric_progression

OpenStudy (anonymous):

Thanks guys; gunna read over it right now.

OpenStudy (paxpolaris):

as the wikipedia article shows you... for 30-60-90 triangles and 45-45-90 triangles you can figure out the ratios of sides just using pythagorenan theorem (and a little geo.)